OpenStudy (anonymous):

the combustion of 6.50 g of butane, C4H10, produces 318.8 Kj of heat. calculate the molar heat of combustion of butane (Kj/mole): (the answer is -2850) I want the method

OpenStudy (anonymous):

If 318.8kJ of energy is released then the enthalpy of combustion of 6.5g of C4H10 is -318.8kJ Find the molar mass of butane: M(C4H10) = 4M(C) + 10M(H) = 4*12 + 10*1 = 58 g/mol \[\Delta _{c}H ^{o}(C _{4}H _{10})=(-318.8) \times \frac{ 58 }{ 6.5 } = 2844.7 kJ/mol \approx 2850kJ/mol\]

OpenStudy (anonymous):

Thank you so much !